OpenStudy (anonymous):

11/12 + 2/3 = 13/15 13/12 19/12 22/36

OpenStudy (isaiah.feynman):

\[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad+bc }{ bd }\]

jimthompson5910 (jim_thompson5910):

hint 2/3 = 8/12 (multiply top and bottom by 4)

jimthompson5910 (jim_thompson5910):

so 11/12 + 2/3 is the same as 11/12 + 8/12

OpenStudy (anonymous):

11/12 +2/3 you need to find a common denominator in this case its simple what can you multiply 3 by to get 12? 4! what you do the bottom of a fraction you must also do to the top so 2/3 x 4 (on the top and the bottom) gives you 8/12 now you can add them 11/12+8/12 gives you 19/12